# A broken debris cascade as a possible source of hot dust emission in transitioning planet-forming disks

### v1

@@ 2.4 Removal from the cascade
Due to the effects of gas described above, particles are removed from the collisional cascade. First of all, particles which are tightly coupled to the gas (with $s<s_{\mathrm{b}}$) move with the gas, removing the relative velocities between them, preventing them from colliding and fragmenting. Secondly, not well-coupled particles are removed from the cascade if the time it takes for them to be damped into the midplane is shorter than the timescale at which they collide. Using Eqs. (13) and (20) we find that this is the case for particles with
@@
In general $s_{\mathrm{cc}}>s_{\mathrm{b}}$, therefore all particles larger than $s_{\mathrm{cc}}$ are involved in the collisional cascade and produce smaller particles through fragmentation. The abundance of small particles is what determines the cloud’s optical depth, since for power-laws with $\gamma$ steeper than $-3$, these dominate the opacity.
@@ 3.1 Size distribution of small particles
In the presence of gas, we therefore have a broken cascade where particles below a certain size no longer collide and fragment. These particles are still produced by the fragmentation of larger particles, but settling is what removes them from the higher regions of the dust cloud, so that they will no longer contribute to the optical depth of the cloud. Taking into account the production and removal of particles, the size distribution of bound particles $f_{\mathrm{b}}(s)$ is given by
@@ 3.2 Production rate of small particles
The production rate of these small ($s\ll s_{\mathrm{b}}$) particles can be derived by considering the result of the fragmentation of a single large, unbound particle. As shown by, among others, Takasawa et al. (2011) and O’Brien & David (2003), the fragments resulting form high-velocity collisions such as those in our system can be assumed to follow a power-law size distribution, i.e.
@@
with $s_{0}$ the initial particle’s size, and $\gamma_{\mathrm{f}}$ between $-3$ and $-6$. Since we are mainly interested in the production of small particles, a more negative value for this power-law would be beneficial. For now we assume the worst case scenario where $\gamma_{\mathrm{f}}=-3$. Using that, assuming the largest fragment to be half the size of the initial particle, mass-conservation requires (O’Brien & David, 2003)
@@
Particles smaller than $s_{\mathrm{cc}}/2$ are produced by all particles in the collisional cascade: from $s_{\mathrm{cc}}$ up to the largest planetesimals of size $s_{\mathrm{pt}}$, and the total production rate of particles with size $\in[s,s+\mathrm{d}s]$ is given by
@@
Using Eq. (20) for the collision time $t_{\mathrm{coll}}$, and Eq. (17) for the size distribution of particles in the collisional cascade $f_{\mathrm{cc}}$, we find
@@
Using $\gamma_{\mathrm{f}}=-3$ and substituting Eq. (21) for $s_{\mathrm{cc}}$, we find
@@ 3.3 Replenishment of large particles
As we can see, the amount of small particles is dependent on the abundance of large, uncoupled particles through $f_{0,\mathrm{cc}}$. To compensate for the constant destruction of these uncoupled particles, we assume that the largest particles in the cascade are constantly being replenished by planetesimals being injected into the system. This can happen in two ways. If there is a population of somewhat larger planetesimals with collisional lifetimes equal to or larger than the lifetime of the disk phase we are talking about (typically less that 10 Myr), the mass injection into the cascade could simply come from collisions of these planetesimals, but would require a sufficient amount of these larger bodies to be present (Löhne et al., 2008; Krivov & Wyatt, 2021). These planetesimals would have to be collisionally excited by planets early on in the disk. An alternative is that smaller planetesimals with shorter collisional lifetimes are injected continuously into relevant orbits by interactions with planets in the disk, for example injecting bodies from further out in the disk by a planet there (Raymond & Izidoro, 2017), or simply by scattering a local planetesimal into a high-inclination orbit (Krijt & Dominik, 2011).
Assuming the size of the largest fragment resulting from the destruction of a particle to be half of that of the initial particle, the mass flux required to keep up the population of the largest particles with $s\in[s_{\mathrm{pt}}/2,s_{\mathrm{pt}}]$, is
@@
Using Eq. (4) for $m$ and Eq. (20) for $t_{\mathrm{coll}}$ we find for the mass flux
@@
Solving for $f_{0,\mathrm{cc}}$ we can express the abundance of (uncoupled) particles in the collisional cascade in terms of the cloud’s dimensions and the mass flux, i.e.
@@
Substituting this for $f_{0,\mathrm{cc}}$ into Eq. (32) we find for the production rate of small particles
$\phi_{\mathrm{prod}}(s)=\frac{9}{2\pi^{2}(\ln{2})^{2}}\frac{1}{\rho_{\mathrm{gas}}^{2}v_{\mathrm{th}}^{2}}\frac{\nu v_{\mathrm{K}}\epsilon_{0}}{V}\dot{M}_{\mathrm{in}}^{2}s^{-3}$ (36)
@@ 3.4 The broken distribution
We can now find the distribution of coupled particles by substituting the equation for the production rate above and Eq. (12) into Eq. (23) for $f_{\mathrm{b}}$, which yields
@@
We can now compare two cases: first of all, in the complete absence of gas, a normal collisional cascade emerges in which planetesimals of size $s_{\mathrm{pt}}$ are supplied and particles below $s_{\mathrm{min}}$ are removed (by radiation pressure, stellar wind or Poynting-Robertson drag (Lagrange et al., 2000; Wyatt, 2008)). The complete size distribution is
@@
Secondly, in the presence of gas, particles smaller than $s_{\mathrm{b}}$ (given by Eq. (7)) are tightly coupled to the gas and follow a distribution with a steeper slope given by Eq. (37). Furthermore, intermediate size particles smaller than $s_{\mathrm{cc}}$ have a reduced abundance, since they are damped into the midplane quicker than they collide, resulting in the distribution given by Eq. (38). The complete size distribution for this case is
@@ 4.1 An opaque cloud
Using this cross-section and a general power-law $f(s)=f_{0}s^{\gamma}$, integration yields
@@
By solving for $f_{0}$ we find the minimum amount of particles needed for a given geometric optical depth
@@
If $\gamma<-3$ (which is the case for small particles in gas as well as for particles in the collisional cascade, see Eqs. 17 and 37) the optical depth is dominated by small particles, and this can be approximated as
@@ 4.2 Mass flux in the presence of gas
which can be simplified by making some rough approximations for the cloud’s structure. First of all, we assume the cloud to be a torus-like shell section of width $\Delta R$ at a distance $R$ to the central star, with an opening half angle $\alpha$ (see Fig. 1). In this way the cloud covers a fraction $\eta\approx\tan{\alpha}$ of the total solid angle from the star. The volume can be approximated as
@@
Furthermore, we approximate the Kepler frequency $\Omega_{\mathrm{K}}$ (and consequentially the Kepler velocity with $v_{\mathrm{K}}=\Omega_{\mathrm{K}}R$) with
@@
for all particles. Furthermore, the cloud’s height is approximately $H=\eta R$, and assuming homogeneity the gas density becomes
@@
with $\Sigma_{\mathrm{gas}}$ the surface density. Finally, given that the largest particles are on orbits with inclinations up to $\alpha$. Therefore, their vertical velocities can get as high as $v_{\mathrm{K}}\tan{\alpha}=\eta v_{\mathrm{K}}$. Since it is mainly these vertical velocity differences that will cause particles to collide, an appropriate number for the average relative velocity between particles is of this order of magnitude, which is why we use $\nu=\eta/4$ (Lissauer & Stewart, 1993, for a typical inclination of 15 degrees). Using all of this and Eq. (2) for $v_{\mathrm{th}}$, Eq. (48) evaluates to
@@ 4.3 Mass flux without gas
The same can be done for the case of a classical debris disk where there is no dynamically relevant gas in the system, and all particles are in the collisional cascade. Equating $f_{0,\mathrm{min}}$ from Eq. (47) with $\gamma=-3.5$ to $f_{0,\mathrm{cc}}$ from Eq. (35), solving for the mass flux yields
@@
Using the same approximations for $V$, $\nu$ and $v_{\mathrm{K}}$ as described above this evaluates to
@@ 4.4 Influence of fragment size distribution
Above $\gamma_{\mathrm{f}}=-3$ was used for the size distribution of the particles resulting from a fragmentation event. This is a very conservative assumption however, as Takasawa et al. (2011) showed that $\gamma_{\mathrm{f}}$ can be as low as $-6$ for collisions at the velocities we are dealing with. As we are mainly interested in the production of small particles, lower values of $\gamma_{\mathrm{f}}$ would be beneficial, and will bring down the required mass flux even further.
By solving the integrals in Eqs. (26), (28) and (43) numerically, we can find $f_{0,\mathrm{min}}$ and $f_{0,\mathrm{b}}$ for different values of $\gamma_{\mathrm{f}}$ as a function of $\dot{M}_{\mathrm{in}}$. Root finding can be used to find the value of $\dot{M}_{\mathrm{in}}$ for which $f_{0,\mathrm{b}}=f_{0,\mathrm{min}}$.
@@ 5.1 Mass flux comparison
Figure 3: The required mass flux $\dot{M}_{\mathrm{in}}$ (in earth masses per year) against the cloud’s geometric optical depth $\tau_{\mathrm{g}}$ (left panel). Both the case where there is no gas (solid line), as well as the one with gas (dashed lines) are given. The latter is plotted for different values of the gas surface density $\Sigma_{\mathrm{gas}}$. In the right panel the required mass flux $\dot{M}_{\mathrm{in}}$ (in earth masses per year) is plotted against the size of the smallest particles $s_{\mathrm{min}}$. Both the case where there is no gas (solid line), as well as the one with gas (dashed lines) are given. The latter is plotted for different values of the gas surface density $\Sigma_{\mathrm{gas}}$.
@@
The equations for the mass flux required for an opaque cloud derived in Sec. 4 show the effect of the presence of gas: using our fiducial parameters (see table 1), the presence of gas decreases the mass flux by a little over an order of magnitude. More importantly, this mass flux can be brought down further by decreasing the gas density, since $\dot{M}_{\mathrm{in}}\propto\Sigma_{\mathrm{gas}}^{0.5}$. This is due to a decreased surface density increasing the time it takes for a planetesimal to be damped into the midplane, causing $s_{\mathrm{cc}}$ to increase and more particles to be involved in the cascade producing small particles.
@@
In the left panel of Fig. 3 we show how the mass flux scales with the optical depth $\tau_{\mathrm{g}}$ for multiple gas (surface) densities (solid lines) and when is no gas at all (dashed black line). First of all, we can see that in the absence of gas, the required mass flux scales with $\tau_{\mathrm{g}}^{2}$ instead of $\tau_{\mathrm{g}}^{0.5}$ in the presence of gas. Or in other words: increasing the mass flux has a stronger effect on the optical depth when there is gas compared to when there is no gas. Increasing the mass flux increases the abundance of particles in the collisional cascade, which decreases their collisional lifetime. In the case where there is no gas present, the lifetime of small particles is their collisional lifetime, and the decline thereof negatively affects the abundance of small particles. The lifetime of small particles in gas, the settling timescale, is unaffected by the mass flux however. Therefore, the effect of increasing the mass flux on the optical depth is stronger when there is gas around. Or, equivalently, increasing the optical depth causes a stronger increase in the required mass flux in the gas-free case.
@@
As a result of this for very low optical depth particles become very sparse, increasing their collision times enough to make the no-gas case preferable. At higher optical depths however settling times are much longer than collision times, and having gas around is preferable. Decreasing the gas surface density by a factor of 1000 has a profound effect and reduces the required mass flux to as little as $10^{-5}M_{\oplus}$ yr-1, an amount that could be sustainable for millions of years during the disk life time.
@@
The right panel of Fig. 3 shows how the required mass flux changes when the smallest particle size deviates from $s_{\mathrm{min}}=0.1$ $\mu$m. Most importantly, having a smaller $s_{\mathrm{min}}$ decreases the mass flux required for a certain optical depth, since smaller particles have a higher surface area to mass ratio, meaning the same mass is more efficient at absorbing light. It should be noted that we are comparing the gas and no gas cases for equal values of $s_{\mathrm{min}}$. When there is no gas around however, small particles are removed almost instantly due to radiation pressure (Backman & Paresce, 1993), whereas these particles would remain in the system when there is gas present. Therefore the effective value of $s_{\mathrm{min}}$ would be higher in a system without gas than it is in one with gas, making the latter more preferable.

### v3

@@ 2.4 Removal from the cascade
Due to the effects of gas described above, particles are removed from the collisional cascade. First of all, particles which are tightly coupled to the gas (with $s<s_{\mathrm{b}}$) move with the gas, removing the relative velocities between them, preventing them from colliding and fragmenting. Secondly, not well coupled particles are removed from the cascade if the time it takes for them to be damped into the midplane is shorter than the timescale at which they collide. Using Eqs. (13) and (20) we find that this is the case for particles with:
@@
In general, $s_{\mathrm{cc}}>s_{\mathrm{b}}$; therefore, all particles larger than $s_{\mathrm{cc}}$ are involved in the collisional cascade and produce smaller particles through fragmentation. The abundance of small particles is what determines the cloud’s optical depth, since for power-laws with $\gamma$ steeper than $-3$, these dominate the opacity.
@@ 3.1 Size distribution of small particles
In the presence of gas, we have now a broken cascade where particles below a certain size no longer collide and fragment. These particles are still produced by the fragmentation of larger particles, but settling is what removes them from the higher regions of the dust cloud, so that they will no longer contribute to the optical depth of the cloud. Taking into account the production and removal of particles, the size distribution of bound particles $f_{\mathrm{b}}(s)$ is given by
@@ 3.2 Production rate of small particles
The production rate of these small ($s\ll s_{\mathrm{b}}$) particles can be derived by considering the result of the fragmentation of a single large, unbound particle. As shown by Takasawa et al. (2011) and O’Brien & David (2003), for instance, the fragments resulting from high-velocity collisions (such as those in our system) can be assumed to follow a power-law size distribution, namely:
@@
with $s_{0}$ the initial particle’s size, and $\gamma_{\mathrm{f}}$ between $-3$ and $-6$. Since we are mainly interested in the production of small particles, a more negative value for this power-law would be beneficial. For now we assume the worst case scenario where $\gamma_{\mathrm{f}}=-3$. Using that, assuming the largest fragment to be half the size of the initial particle, mass-conservation requires (O’Brien & David, 2003):
@@
Particles smaller than $s_{\mathrm{cc}}/2$ are produced by all particles in the collisional cascade: from $s_{\mathrm{cc}}$ up to the largest planetesimals of size $s_{\mathrm{pt}}$, and the total production rate of particles with sizes of $\in[s,s+\mathrm{d}s]$ is given by:
@@
Using Eq. (20) for the collision time $t_{\mathrm{coll}}$, and Eq. (17) for the size distribution of particles in the collisional cascade $f_{\mathrm{cc}}$, we find:
@@
Using $\gamma_{\mathrm{f}}=-3$ and substituting Eq. (21) for $s_{\mathrm{cc}}$, we find:
@@ 3.3 Replenishment of large particles
As we can see, the amount of small particles is dependent on the abundance of large, uncoupled particles through $f_{0,\mathrm{cc}}$. To compensate for the constant destruction of these uncoupled particles, we assume that the largest particles in the cascade are constantly being replenished by planetesimals being injected into the system. This can happen in two ways. If there is a population of somewhat larger planetesimals with collisional lifetimes equal to or larger than the lifetime of the disk phase we are talking about (typically less that 10 Myr), the mass injection into the cascade could simply come from collisions of these planetesimals, but would require a sufficient amount of these larger bodies to be present (Löhne et al., 2008; Krivov & Wyatt, 2021). These planetesimals would have to be collisionally excited by planets early on in the disk. An alternative is that smaller planetesimals with shorter collisional lifetimes are injected continuously into relevant orbits by interactions with planets in the disk, for example, by injecting bodies from further out in the disk by a planet there (Raymond & Izidoro, 2017) or simply by scattering a local planetesimal into a high-inclination orbit (Krijt & Dominik, 2011).
Assuming the size of the largest fragment resulting from the destruction of a particle to be half of that of the initial particle, the mass flux required to keep up the population of the largest particles with $s\in[s_{\mathrm{pt}}/2,s_{\mathrm{pt}}]$, is:
@@
Using Eq. (4) for $m$ and Eq. (20) for $t_{\mathrm{coll}}$, for the mass flux we find:
@@
Solving for $f_{0,\mathrm{cc}}$ we can express the abundance of (uncoupled) particles in the collisional cascade in terms of the cloud’s dimensions and the mass flux, namely,
@@
Substituting this for $f_{0,\mathrm{cc}}$ into Eq. (32), we find the production rate of small particles to be:
$\phi_{\mathrm{prod}}(s)=\frac{9}{2\pi^{2}(\ln{2})^{2}}\frac{1}{\rho_{\mathrm{gas}}^{2}v_{\mathrm{th}}^{2}}\frac{\nu v_{\mathrm{K}}\epsilon_{0}}{V}\dot{M}_{\mathrm{in}}^{2}s^{-3}.$ (36)
@@ 3.4 The broken distribution
We can now find the distribution of coupled particles by substituting the equation for the production rate above and Eq. (12) into Eq. (23) for $f_{\mathrm{b}}$, which yields:
@@
We can now compare two cases: first of all, in the complete absence of gas, a normal collisional cascade emerges in which planetesimals of size, $s_{\mathrm{pt}}$, are supplied and particles below $s_{\mathrm{min}}$ are removed (by radiation pressure, stellar wind or Poynting-Robertson drag (Lagrange et al., 2000; Wyatt, 2008)). The complete size distribution is
@@
Secondly, in the presence of gas, particles smaller than $s_{\mathrm{b}}$ (given by Eq. (7)) are tightly coupled to the gas and follow a distribution with a steeper slope given by Eq. (37). Furthermore, intermediate size particles smaller than $s_{\mathrm{cc}}$ have a reduced abundance, since they are damped into the midplane quicker than they collide, resulting in the distribution given by Eq. (38). The complete size distribution for this case is:
@@ 4.1 An opaque cloud
Using this cross-section and a general power-law $f(s)=f_{0}s^{\gamma}$, integration yields:
@@
By solving for $f_{0}$ we find the minimum amount of particles needed for a given geometric optical depth as:
@@
If $\gamma<-3$ (which is the case for small particles in gas as well as for particles in the collisional cascade, see Eqs. 17 and 37) the optical depth is dominated by small particles, and this can be approximated as:
@@ 4.2 Mass flux in the presence of gas
which can be simplified by making some rough approximations for the cloud’s structure. First of all, we assume the cloud to be a torus-like shell section of width $\Delta R$ at a distance $R$ to the central star, with an opening half angle $\alpha$ (see Fig. 1). In this way the cloud covers a fraction of $\eta\approx\tan{\alpha}$ of the total solid angle from the star. The volume can be approximated as:
@@
Furthermore, we approximate the Kepler frequency $\Omega_{\mathrm{K}}$ (and consequentially the Kepler velocity with $v_{\mathrm{K}}=\Omega_{\mathrm{K}}R$) with:
@@
for all particles. Furthermore, the cloud’s height is approximately $H=\eta R$, and assuming homogeneity the gas density becomes:
@@
with $\Sigma_{\mathrm{gas}}$ the surface density. Finally, given that the largest particles are on orbits with inclinations up to $\alpha$. Therefore, their vertical velocities can get as high as $v_{\mathrm{K}}\tan{\alpha}=\eta v_{\mathrm{K}}$. Since it is mainly these vertical velocity differences that will cause particles to collide, an appropriate number for the average relative velocity between particles is of this order of magnitude, which is why we use $\nu=\eta/4$ (Lissauer & Stewart, 1993, for a typical inclination of 15 degrees). Using all of this and Eq. (2) for $v_{\mathrm{th}}$, we can evaluate Eq. (48) to:
@@ 4.3 Mass flux without gas
The same can be done for the case of a classical debris disk where there is no dynamically relevant gas in the system, and all particles are in the collisional cascade. Equating $f_{0,\mathrm{min}}$ from Eq. (47) with $\gamma=-3.5$ to $f_{0,\mathrm{cc}}$ from Eq. (35), solving for the mass flux yields:
@@
Using the same approximations for $V$, $\nu$ and $v_{\mathrm{K}}$ as described above, this can be expressed as:
@@ 4.4 Influence of fragment size distribution
Above, $\gamma_{\mathrm{f}}=-3$ is used for the size distribution of the particles resulting from a fragmentation event. This is a very conservative assumption, however, as Takasawa et al. (2011) showed that $\gamma_{\mathrm{f}}$ can be as low as $-6$ for collisions at the velocities we are dealing with. As we are mainly interested in the production of small particles, lower values of $\gamma_{\mathrm{f}}$ would be beneficial and would bring down the required mass flux even further. By solving the integrals in Eqs. (26), (28) and (43) numerically, we can find $f_{0,\mathrm{min}}$ and $f_{0,\mathrm{b}}$ for different values of $\gamma_{\mathrm{f}}$ as a function of $\dot{M}_{\mathrm{in}}$. Root finding can be used to find the value of $\dot{M}_{\mathrm{in}}$ for which $f_{0,\mathrm{b}}=f_{0,\mathrm{min}}$.
@@ 5.1 Mass flux comparison
Figure 3: Required mass flux $\dot{M}_{\mathrm{in}}$ (in earth masses per year) against the cloud’s geometric optical depth $\tau_{\mathrm{g}}$ (left panel). Both the case where there is no gas (solid line), as well as the one with gas (dashed lines) are given. The latter is plotted for different values of the gas surface density, $\Sigma_{\mathrm{gas}}$. In the right panel the required mass flux, $\dot{M}_{\mathrm{in}}$ (in earth masses per year), is plotted against the size of the smallest particles, $s_{\mathrm{min}}$. Both the case where there is no gas (solid line), as well as the one with gas (dashed lines) are given. The latter is plotted for different values of the gas surface density, $\Sigma_{\mathrm{gas}}$.
@@
The equations for the mass flux required for an opaque cloud (derived in Sect. 4) show the effect of the presence of gas: using our fiducial parameters (see table 1), the presence of gas decreases the mass flux by a little over an order of magnitude. More importantly, this mass flux can be brought down further by decreasing the gas density, since $\dot{M}_{\mathrm{in}}\propto\Sigma_{\mathrm{gas}}^{0.5}$. This is due to a decreased surface density increasing the time it takes for a planetesimal to be damped into the midplane, causing $s_{\mathrm{cc}}$ to increase and more particles to be involved in the cascade producing small particles.
@@
In the left panel of Fig. 3, we show how the mass flux scales with the optical depth, $\tau_{\mathrm{g}}$, for multiple gas (surface) densities (solid lines) and when is no gas at all (dashed black line). First of all, we can see that in the absence of gas, the required mass flux scales with $\tau_{\mathrm{g}}^{2}$ instead of $\tau_{\mathrm{g}}^{0.5}$ in the presence of gas. In other words: increasing the mass flux has a stronger effect on the optical depth when there is gas compared to when there is no gas. Increasing the mass flux increases the abundance of particles in the collisional cascade, which decreases their collisional lifetime. In the case where there is no gas present, the lifetime of small particles is their collisional lifetime, and the decline thereof negatively affects the abundance of small particles. The lifetime of small particles in gas, the settling timescale, is unaffected by the mass flux however. Therefore, the effect of increasing the mass flux on the optical depth is stronger when there is gas around; or, equivalently, increasing the optical depth causes a stronger increase in the required mass flux in the gas-free case.
@@
As a result of this for very low optical depth particles become very sparse, increasing their collision times enough to make the no-gas case preferable. At higher optical depths however settling times are much longer than collision times, and having gas around is preferable. Decreasing the gas surface density by a factor of 1000 has a profound effect and reduces the required mass flux to as little as $10^{-5}M_{\oplus}$ yr-1, an amount that could be sustainable for millions of years during the disk’s lifetime.
@@
The right panel of Fig. 3 shows how the required mass flux changes when the smallest particle size deviates from $s_{\mathrm{min}}=0.1$ $\mu$m. Most importantly, having a smaller $s_{\mathrm{min}}$ decreases the mass flux required for a certain optical depth, since smaller particles have a higher surface area to mass ratio, meaning the same mass is more efficient at absorbing light. It should be noted that we are comparing the gas and no gas cases for equal values of $s_{\mathrm{min}}$. When there is no gas present, however, small particles are removed almost instantly due to radiation pressure (Backman & Paresce, 1993); whereas these particles would remain in the system when there is gas present. Therefore, the effective value of $s_{\mathrm{min}}$ would be higher in a system without gas than it is in one with gas, making the latter more preferable.